A 1200 kg car is propelled down a level road by a force of 300N. There are 500N of friction and 1000N of Air resistence. What is the acceleration of the car? What does the car weigh? .
Answer #1

a)

From newton's second law

-FApplied+FFriction+FAir =ma

-300+500+1000=1200a

a=1200/1200 =1 m/s2

b)

Weight of car

W=mg =1200*9.8 =11760 N